OpenStudy (anonymous):

Solve . A. 3 ≤ x ≤ 52 B. x ≥ 3 C. x ≥ 52 D. x ≤ 52

OpenStudy (anonymous):

ops solve.\[\sqrt{x-3}\ge7\]

OpenStudy (anonymous):

i think its B not sure

OpenStudy (jdoe0001):

\(\bf \sqrt{x-3}\ge 7\quad \textit{square both sides}\\ \quad \\ (\sqrt{x-3})^2\ge 7^2\implies \sqrt{(x-3)}^2\ge 7^2\implies x-3 \ge 7^2\) I gather you can get it from there